there is a license plate with 7 spots. the first 4 are capital letters and the last 3 are digits. what is the probability that the first 3 letters will spell "wow"?
Answer by stanbon(75887)

there is a license plate with 7 spots. the first 4 are capital letters and the last 3 are digits. what is the probability that the first 3 letters will spell "wow"?
---
You are selecting with replacement so the selections are independent = .
P(wow) = (1/26)^3 = (1/26)^3 = 0.00005690,,
---
Cheers,
Stan H.